OpenStudy (anonymous):

anybody know whats x + 8 =-10?

OpenStudy (anonymous):

Ok. You want to get x by itself on both sides, so you need to get rid of that 8. How do you think you could do that?

OpenStudy (anonymous):

I'm thinking of a number, and 8 more than this number is -10.

OpenStudy (anonymous):

add 8 to both sides i think lol

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Subtract 8 from both sides. +8 - 8 = 0

OpenStudy (anonymous):

-10 - 8 =

OpenStudy (anonymous):

okay so a negative minus a positive is a postive?

OpenStudy (anonymous):

Nope, a positive minus a negative is a positive. A negative minus a positive is a negative.

OpenStudy (anonymous):

What is 8 less than -10?

OpenStudy (anonymous):

so -10 - 8=-18

OpenStudy (anonymous):

Ok, if you like his solution, click the "good answer" button by his answer.

OpenStudy (anonymous):

Ok, so we have: \(x + 8 =-10\) \[\implies x + 8 - 8 = -10 - 8\]\[\implies x + 0 = -18\]\[\implies x = -18\]

OpenStudy (anonymous):

! someone deleted my reply. :(

OpenStudy (anonymous):

x+8=-10 -8=-8 ___________ x=-18 the positive 8 x's out the negative 8. hope this helps! :)

OpenStudy (anonymous):

thanks i got that one wrong but i knew it ugh..thank you